A shopping centre car park has spaces for 8 buses, 160 cars and 12 motorbikes. If all vehicles have an equal chance of leaving at any time, find, expressing your answer as a proper fraction in its simplest form
1
Expert's answer
2019-10-21T09:39:37-0400

Total number of vehicles: "N=8+160+12=180."

Probability of bus: "\frac{8}{180}=\frac{2}{45}."

Probability of car: "\frac{160}{180}=\frac{8}{9}."

Probability of motorbike: "\frac{12}{180}=\frac{1}{15}."